OpenStudy (anonymous):

An inlet pipe can fill a tank in 10 minutes. A drain can empty the tank in 12 minutes. If the tank is empty and both pipe and drain are open, how long will it be before the tank overflows?

OpenStudy (amistre64):

Total Volume = volume in - volume out each minute. Tv = 1/10 - 1/12 per minute. Tv = (12 - 10)/120 = 2/120 = 1/60 per minute. After 60 minutes, the tank should be overflowing.

OpenStudy (anonymous):

Excellent answer amistre64! Wow.